OpenStudy (anonymous):

There is a 25ft ladder leaned 7ft away from a building if the top of the ladder slips down 4ft, how many feet will the bottom slide out?

OpenStudy (whpalmer4):

Draw a picture: 

OpenStudy (anonymous):

4 ft?

OpenStudy (whpalmer4):

 By Pythagorean theorem, we have \[7^2 + x^2 = 25^2\] The ladder stays the same length when the height along the wall goes from \(x\) to \(x-4\). That means that the bottom leg is now (7+y) and\[(x-4)^2 + (7+y)^2 = 25^2\]

OpenStudy (whpalmer4):

If you like, you can solve for the value of \(x\) before attempting to find \(y\).

OpenStudy (whpalmer4):

In fact, you probably should :-)

OpenStudy (anonymous):

oh my bad its 8 ft

OpenStudy (whpalmer4):

Yes, that's correct.